Two trains leave a station at the same time are going in the same direction. One travels at 40 mi/h and the other travels at 55 mi/h. In how many hours will they be 72 miles apart?
Answer by chrisrao(11)

S = (V1-V2)T
72 = (55-40)T
T = 72/15 = 4.8 hours